Determine the kth percentile of the distribution defined by the following density function. f (x)= x o SxSVE O elsewhere
Answer #1

We want to find 'x' such that P(X<x)=\frac{k}{100}

\\\int_{0}^{x}(x)dx=\frac{k}{100}\\ \\

[\frac{x^2}{2}]_{0}^{x}=\frac{k}{100}\\ \\

[\frac{x^2}{2}-0]=\frac{k}{100}\\ \\

\Rightarrow x=\sqrt{\frac{k}{50}}

which is the kth percentile.